# Koopman for stochastic dynamics: error bounds for kernel extended dynamic mode decomposition

### v1

@@ 4.2. Probabilistic approximation step: estimating the expected value
$\mathbb{P}\big(\|g-g^{\mathrm{MC}}\|_{2}\geq\varepsilon\|f\|_{H})\leq 2\exp\left(-\frac{m\varepsilon^{2}}{8d\|k\|_{\infty}}\right).$ (43)
By the $\ell^{1}$-$\ell^{2}$-inequality, the deterministic second term is smaller than the one in Theorem 4.6, i.e., $\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{2}\leq\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}$. We refer the reader to Appendix A for detailed calculations. In total, with probability at least $1-2\exp\left(-\frac{m\varepsilon^{2}}{8d\|k\|_{\infty}}\right)$, the bound
@@
$\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}\leq\|k\|_{\infty}^{1/2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$
@@
From the positive definiteness of the kernel $k$, we get the positive semidefiniteness of the augmented kernel matrix $\mathbf{K}_{\mathcal{X}\cup\{x\}}=\begin{pmatrix}\mathbf{K}_{\mathcal{X}}&\mathbf{k}_{\mathcal{X}}\\ \mathbf{k}_{\mathcal{X}}^{\top}&k(x,x)\end{pmatrix}$ with positive semidefinite (thus, non-negative) Schur complement $k(x,x)-\mathbf{k}_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}\geq 0$. With $\lambda_{\mathrm{max}}(\mathbf{K}_{\mathcal{X}}^{-1})=\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}})^{-1}$, where $\lambda_{\max}$ and $\lambda_{\min}$ are the largest and smallest eigenvalue of the respective matrices, we get
$\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}^{2}\leq k(x,x)\cdot\frac{d}{\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}})}.$
The term in the denominator usually decays rapidly as the condition of the kernel matrix gets worse. If our kernel is given by a radial basis function (RBF) $B:\mathbb{R}^{n}\to\mathbb{R}$, $k(x,x)=B(0)$ is constant, so
@@
For example, for the Wendland kernel $k_{n,l}$ with RBF $B_{n,l}$, we have $B_{n,l}(0)=\frac{1}{2^{l}}\binom{\lfloor n/2\rfloor+2l+1}{l}$, while for Matérn kernels $k_{\nu}$ with RBF $\mathcal{M}_{\nu}$, we have $\mathcal{M}_{\nu}(0)=1$.
@@
Let the assumptions of of Theorem 4.6 hold. For $f\in\mathbb{H}$, the regularized approximant $\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}$ from (44) satisfies
$\|(\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}-\widehat{\mathcal{K}}_{\lambda})f\|_{\infty}\,\leq\,\varepsilon\cdot\|k\|_{\infty}^{1/2}\cdot\sqrt{\frac{d}{\lambda}}\|f\|_{H}$ (45)
with probability at least $1-2de^{-m\varepsilon^{2}/(2\|k\|_{\infty})}$.
@@
$\|(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}^{2}\leq k(x,x)\frac{d}{\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}}+\lambda I)}\leq k(x,x)\frac{d}{\lambda}.$
@@
$\|g_{\lambda}-g_{\lambda}^{\mathrm{MC}}\|_{\infty}=\max_{x_{j}\in\mathcal{X}}\left|\mathbb{E}[(R_{\mathcal{X}}^{\lambda}f)(Y)|X=x_{j}]-\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}(R_{\mathcal{X}}^{\lambda}f)(y_{l,j})\right|$
with the operator $R_{\mathcal{X}}^{\lambda}$ from (37). By [BoldPhil24, Proposition B.1], that “projection-like” operator satisfies $\|R_{\mathcal{X}}^{\lambda}\|_{\mathbb{H}\to\mathbb{H}}\leq\|S_{\mathcal{X}}\|_{\mathbb{H}\to\mathbb{H}}=1$. Thus, the arguments from the proof of Theorem 4.6 carry over, and for every $f\in\mathbb{H}$, we get
$\mathbb{P}\left(\|g_{\lambda}-g_{\lambda}^{\mathrm{MC}}\|_{\infty}\geq\varepsilon\|f\|_{\mathbb{H}}\right)\leq 2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right).$
@@
The appeal of Corollary 4.9 lies in the more explicit bound in terms of the number $d$ of data points in $\mathcal{X}$. The regularization parameter $\lambda$ puts a “floor” on the conditioning of the kernel matrix $\mathbf{K}_{\mathcal{X}}$, making the tradeoff between deterministic regularization bias and kernel matrix conditioning evident.
In the final subsection, we combine the deterministic and Monte Carlo error bounds to get a bound on the full $L^{\infty}$-error of the approximations $\widehat{\mathcal{K}}^{\mathrm{MC}}$ and $\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}$.
@@
Let $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b},C_{b})$. Furthermore, for $\mathcal{X}=\{x_{1},\ldots,x_{d}\}$, let $m_{j}=m\in\mathbb{N}$ i.i.d. samples from every $\rho_{x_{j}}$, $j\in\{1,\dots,d\}$. Then, for every $\varepsilon>0$ and for all $f\in\mathbb{H}$, $\|f\|_{\mathbb{H}}\leq 1$, we have the pointwise error bound
$\|(\mathcal{K}-\widehat{\mathcal{K}}^{\mathrm{MC}})f\|_{\infty}\leq\big[1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}\big]\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}+\varepsilon\cdot\|k\|_{\infty}^{1/2}\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right)$.
@@
$\|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f\|_{\infty}\,\leq\,\varepsilon\cdot\|k\|_{\infty}^{1/2}\cdot\big(\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\big)^{1/2}\|f\|_{H}$
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right)$. With $\|f\|_{\mathbb{H}}\leq 1$, those two inequalities yield the claim. ∎
The projection error $\mathcal{I}-S_{\mathcal{X}}$ and the expression involving the kernel matrix $\mathbf{K}_{\mathcal{X}}^{-1}$ warrant a closer inspection in certain special cases. For Sobolev spaces and Wendland RKHS, the projection error in Theorem 2.8 is bounded in terms of the fill distance $h_{\mathcal{X}}$.
@@
Let $l\in\mathbb{N}$, define $\sigma_{n,l}=\frac{n+1}{2}+l$ and denote by $\mathbb{H}$ the Sobolev space $H^{\sigma_{n,l}}$ induced by the Wendland kernel $k_{n,l}$ as RKHS. Assume that $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b},C_{b})$. Then, for every $\varepsilon>0$ and all $f\in\mathbb{H}$ with $\|f\|_{\mathbb{H}}\leq 1$, we have the pointwise error bound
@@
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2C_{2}^{2}}\right)$, with constants $C_{1},C_{2}>0$ only depending on the kernel $k_{n,l}$.
@@
$\|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f\|_{\infty}\,\leq\,\varepsilon\cdot\|k_{n,l}\|_{\infty}^{1/2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}\|f\|_{H}$
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k_{n,l}\|_{\infty}}\right)$. Setting $C_{2}=\|k_{n,l}\|_{\infty}^{1/2}$ and taking both inequalities for $\|f\|_{\mathbb{H}}\leq 1$ yields the claim. ∎
Note that Wendland kernels are defined via $k_{n,l}(x,y)=B_{n,l}(\|x-y\|_{2})$ with the radial basis function $B_{n,l}:[0,\infty)\to\mathbb{R}$, see Appendix B. It can be shown that $\|k_{n,l}\|_{\infty}=\sup_{x\in\mathbb{R}^{n}}|k_{n,l}(x,x)|=B_{n,l}(0)=p_{n,l}(0)$ (see, e.g., [PhilSchaWort2025]), where $p_{n,l}$ is the recursively defined polynomial underlying the Wendland kernel [Wend04]. Thus, we get the explicitly computable expression $C_{2}=\sqrt{p_{n,l}(0)}$ for the constant from (21) that depends on dimension $n$ and smoothness $l$. A similar bound as in Theorem 2.8 is easily proven for Matérn kernels, which we state in Corollary C.2 in Appendix C.
To make use of the error bounds from Theorem 2.8 in practice, one would first choose a large enough $d$ to ensure that the fill distance $h_{\mathcal{X}}$ controlling the deterministic projection error is sufficiently small. This might result in a badly conditioned kernel matrix $\mathbf{K}_{\mathcal{X}}$, making the expression $\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$ large. Thus, to control the probabilistic error, one must choose a small enough $\varepsilon>0$ that compensates for the conditioning. Lastly, we want our error bound to hold with probability close to one, so the amount of Monte Carlo samples $m$ for the entries of $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ has to guarantee that $2d\exp(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}})$ is close to zero. Since the factor involving the inverse kernel matrix $\mathbf{K}_{\mathcal{X}}^{-1}$ depends on the size $d$ of data set $\mathcal{X}$, regularization as indicated in Corollary 4.9 prevents this expression from blowing up with $d$. However, regularization introduces a bias to the deterministic error, which is why the final result of this section explicitly juxtaposes deterministic and probabilistic error for the regularized approximant $\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}$.
@@
$\|(\mathcal{K}-\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}})f\|_{\infty}\leq\left(\|1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}\right)\,\|\mathcal{I}-R_{\mathcal{X}}^{\lambda}\|_{\mathbb{H}\to C_{b}}\,+\varepsilon\cdot\|k\|_{\infty}^{1/2}\cdot\sqrt{\frac{d}{\lambda}}$ (46)
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right)$. Specifically, for the RKHS $\mathbb{H}=H^{\sigma_{n,l}}$ induced by the Wendland kernel $k_{n,l}$, where $\sigma_{n,l}=\frac{n+1}{2}+l$, we have the pointwise error bound
$\|(\mathcal{K}-\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}})f\|_{\infty}\leq C_{1}\left(h_{\mathcal{X}}^{l+1/2}+\lambda^{1/2}\right)+\varepsilon\cdot C_{2}\cdot\sqrt{\frac{d}{\lambda}}$ (47)
for every $\varepsilon>0$ and all $f\in H^{\sigma_{n,l}}$ with $\|f\|_{H^{\sigma_{n,l}}}\leq 1$ with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2C_{2}^{2}}\right)$ and constants $C_{1},C_{2}>0$ only depending on $k_{n,l}$.
@@
The second claim combines the second part of Proposition 4.5 for the deterministic error bound with setting $C_{2}=\|k_{n,l}\|_{\infty}^{1/2}$, obtaining (47). ∎

### v2

@@ 4.2. Probabilistic approximation step: estimating the expected value
$\mathbb{P}\big(\|g-g^{\mathrm{MC}}\|_{2}\geq\varepsilon\|f\|_{H})\leq 2\exp\left(-\frac{m\varepsilon^{2}}{8d\|\text{\sf{k}}\|_{\infty}}\right).$ (44)
By the $\ell^{1}$-$\ell^{2}$-inequality, the deterministic second term is smaller than the one in Theorem 4.6, i.e., $\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{2}\leq\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}$. We refer the reader to Appendix A for detailed calculations. In total, with probability at least $1-2\exp\left(-\frac{m\varepsilon^{2}}{8d\|\text{\sf{k}}\|_{\infty}}\right)$, the bound
@@
$\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}\leq\|\text{\sf{k}}\|_{\infty}^{1/2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$
@@
From the positive definiteness of the kernel k, we get the positive semidefiniteness of the augmented kernel matrix $\mathbf{K}_{\mathcal{X}\cup\{x\}}=\left(\begin{smallmatrix}\mathbf{K}_{\mathcal{X}}&\mathbf{k}_{\mathcal{X}}\\ \mathbf{k}_{\mathcal{X}}^{\top}&\text{\sf{k}}(x,x)\end{smallmatrix}\right)$ with positive semidefinite (thus, non-negative) Schur complement $\text{\sf{k}}(x,x)-\mathbf{k}_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}\geq 0$. With $\lambda_{\mathrm{max}}(\mathbf{K}_{\mathcal{X}}^{-1})=\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}})^{-1}$, where $\lambda_{\max}$ and $\lambda_{\min}$ are the largest and smallest eigenvalue of the respective matrices, we get
$\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}^{2}\leq\text{\sf{k}}(x,x)\cdot\frac{d}{\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}})}.$
The term in the denominator usually decays rapidly as the condition of the kernel matrix gets worse. If our kernel is given by a radial basis function (RBF) $B:\mathbb{R}^{n}\to\mathbb{R}$, $\text{\sf{k}}(x,x)=B(0)$ is constant, so
@@
For example, for the Wendland kernel $\text{\sf{k}}_{n,l}$ with RBF $B_{n,l}$, we have $B_{n,l}(0)=\frac{1}{2^{l}}\binom{\lfloor n/2\rfloor+2l+1}{l}$, while for Matérn kernels $\text{\sf{k}}_{\nu}$ with RBF $\mathcal{M}_{\nu}$, we have $\mathcal{M}_{\nu}(0)=1$.
@@
Let the assumptions of Theorem 4.6 hold. For $f\in\mathbb{H}$, the regularized approximant $\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}$ from (45) satisfies
$\left\|(\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}-\widehat{\mathcal{K}}_{\lambda})f\right\|_{\infty}\,\leq\,\varepsilon\cdot\|\text{\sf{k}}\|_{\infty}^{1/2}\cdot\sqrt{\frac{d}{\lambda}}\,\|f\|_{\mathbb{H}}$ (46)
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}}\right)$.
@@
$\|(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}^{2}\leq\text{\sf{k}}(x,x)\frac{d}{\lambda_{\mathrm{min}}(\mathbf{K}_{\mathcal{X}}+\lambda I)}\leq\text{\sf{k}}(x,x)\frac{d}{\lambda}.$
@@
$\|g_{\lambda}-g_{\lambda}^{\mathrm{MC}}\|_{\infty}=\max_{x_{j}\in\mathcal{X}}\left|\mathbb{E}[(R_{\mathcal{X}}^{\lambda}f)(Y)\mid X=x_{j}]-\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}(R_{\mathcal{X}}^{\lambda}f)(y_{l,j})\right|$
with the operator $R_{\mathcal{X}}^{\lambda}$ from (38). By [7, Proposition B.1], that “projection-like” operator satisfies $\|R_{\mathcal{X}}^{\lambda}\|_{\mathbb{H}\to\mathbb{H}}\leq\|S_{\mathcal{X}}\|_{\mathbb{H}\to\mathbb{H}}=1$. Thus, the arguments from the proof of Theorem 4.6 carry over, and for every $f\in\mathbb{H}$, we get
$\mathbb{P}\left(\|g_{\lambda}-g_{\lambda}^{\mathrm{MC}}\|_{\infty}\geq\varepsilon\|f\|_{\mathbb{H}}\right)\leq 2d\exp\left(-\frac{m\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}}\right).$
@@
The appeal of Corollary 4.9 lies in the more explicit bound in terms of the number $d$ of data points in $\mathcal{X}$. The regularization parameter $\lambda$ puts a “floor” on the smallest eigenvalue of the kernel matrix $\mathbf{K}_{\mathcal{X}}$, making the tradeoff between deterministic regularization bias and kernel matrix conditioning evident.
In the final subsection, we combine the deterministic and Monte Carlo error bounds to get a bound on the full $L^{\infty}$-error of the approximations $\widehat{\mathcal{K}}^{\mathrm{MC}}$ and $\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}$, proving the remaining main results from Section 2.
@@
Let $\mathbb{H}$ be an RKHS of functions over $\Omega$, and assume $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b}(\Omega),C_{b}(\Omega))$. Furthermore, for $\mathcal{X}=\{x_{1},\ldots,x_{d}\}$, let $m_{j}=m\in\mathbb{N}$ i.i.d. samples from every $\rho_{x_{j}}$, $j\in\{1,\dots,d\}$, be given. Then, for every $\varepsilon>0$ satisfying $m\geq\frac{2\|\text{\sf{k}}\|_{\infty}}{\varepsilon^{2}}\log(2d)$ and for all $f\in\mathbb{H}$, $\|f\|_{\mathbb{H}}\leq 1$, we have the pointwise error bound
$\big\|\big(\mathcal{K}-\widehat{\mathcal{K}}^{\mathrm{MC}}\big)f\big\|_{\infty}\leq\big[1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}\big]\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}+\varepsilon\cdot\|\text{\sf{k}}\|_{\infty}^{1/2}\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}}\right)$, where $\mathbbm{1}=\{v\in\mathbb{R}^{d}:|v_{i}|=1\}$.
@@
$\|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f\|_{\infty}\,\leq\,\varepsilon\cdot\|\text{\sf{k}}\|_{\infty}^{1/2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}\,\|f\|_{H}$
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}}\right)$. With $\|f\|_{\mathbb{H}}\leq 1$, those two inequalities yield the claim. ∎
The projection error $\mathcal{I}-S_{\mathcal{X}}$ and the expression involving the kernel matrix $\mathbf{K}_{\mathcal{X}}^{-1}$ warrant a closer inspection in certain special cases. For Sobolev spaces as Wendland RKHS, the projection error in Theorem 2.8 is bounded in terms of the fill distance $h_{\mathcal{X}}$.
@@
Let $l\in\mathbb{N}$, define $\sigma_{n,l}=\frac{n+1}{2}+l$ and denote by $\mathbb{H}$ the Sobolev space $H^{\sigma_{n,l}}(\Omega)$ induced by the Wendland kernel $\text{\sf{k}}_{n,l}$ as RKHS. Assume that $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b}(\Omega),C_{b}(\Omega))$. Furthermore, for $\mathcal{X}=\{x_{1},\ldots,x_{d}\}\subset\Omega$, let $m_{j}=m\in\mathbb{N}$ i.i.d. samples from every $\rho_{x_{j}}$, $j\in\{1,\dots,d\}$, be given. Then, for every $\varepsilon>0$ satisfying $m\geq\frac{2C_{2}^{2}}{\varepsilon^{2}}\log(2d)$ and all $f\in\mathbb{H}$ with $\|f\|_{\mathbb{H}}\leq 1$, we have the pointwise error bound
@@
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2C_{2}^{2}}\right)$, where $C_{1},C_{2}>0$ are constants only depending on the kernel $\text{\sf{k}}_{n,l}$, and $\mathbbm{1}=\{v\in\mathbb{R}^{d}:|v_{i}|=1\}$.
@@
$\|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f\|_{\infty}\,\leq\,\varepsilon\cdot\|\text{\sf{k}}_{n,l}\|_{\infty}^{1/2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}\|f\|_{H}$
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|\text{\sf{k}}_{n,l}\|_{\infty}}\right)$. Setting $C_{2}=\|\text{\sf{k}}_{n,l}\|_{\infty}^{1/2}$ and taking both inequalities for $\|f\|_{\mathbb{H}}\leq 1$ yields the claim. ∎
Note that Wendland kernels are defined via $\text{\sf{k}}_{n,l}(x,y)=B_{n,l}(\|x-y\|_{2})$ with the radial basis function $B_{n,l}:[0,\infty)\to\mathbb{R}$, see Appendix B. It can be shown that $\|\text{\sf{k}}_{n,l}\|_{\infty}=\sup_{x\in\mathbb{R}^{n}}|\text{\sf{k}}_{n,l}(x,x)|=B_{n,l}(0)=p_{n,l}(0)$ (see, e.g., [41]), where $p_{n,l}$ is the recursively defined polynomial underlying the Wendland kernel [50]. Thus, we get the explicitly computable expression $C_{2}=\sqrt{p_{n,l}(0)}$ for the constant from (22) that depends on dimension $n$ and smoothness $l$. A similar bound as in Theorem 2.8 is easily proven for Matérn kernels, which we state in Corollary C.2 in Appendix C.
To make use of the error bounds from Theorem 2.8 in practice, one would first choose a large enough $d$ to ensure that the fill distance $h_{\mathcal{X}}$ controlling the deterministic projection error is sufficiently small. This might result in a badly conditioned kernel matrix $\mathbf{K}_{\mathcal{X}}$, making the expression $\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}$ large. Thus, to control the probabilistic error, one must choose a small enough $\varepsilon>0$ that compensates for the conditioning. Lastly, we want our error bound to hold with probability close to one, so the amount of Monte Carlo samples $m$ for the entries of $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ has to guarantee that $2d\exp(-\frac{m\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}})$ is close to zero. Since the factor involving the inverse kernel matrix $\mathbf{K}_{\mathcal{X}}^{-1}$ depends on the size $d$ of data set $\mathcal{X}$, regularization as indicated in Corollary 4.9 prevents this expression from blowing up with $d$. However, regularization introduces a bias to the deterministic error, which is why the final result of this section explicitly juxtaposes deterministic and probabilistic error for the regularized approximant $\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}$.
@@
$\big\|\big(\mathcal{K}-\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}\big)f\big\|_{\infty}\leq\left(\|1+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}\right)\,\|\mathcal{I}-R_{\mathcal{X}}^{\lambda}\|_{\mathbb{H}\to C_{b}}\,+\varepsilon\cdot\|\text{\sf{k}}\|_{\infty}^{1/2}\cdot\sqrt{\frac{d}{\lambda}}$ (47)
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}}\right)$. Specifically, for the RKHS $\mathbb{H}=H^{\sigma_{n,l}}(\Omega)$ induced by the Wendland kernel $\text{\sf{k}}_{n,l}$, where $\sigma_{n,l}=\frac{n+1}{2}+l$, we have the pointwise error bound
$\big\|\big(\mathcal{K}-\widehat{\mathcal{K}}_{\lambda}^{\mathrm{MC}}\big)f\big\|_{\infty}\leq C_{1}\left(h_{\mathcal{X}}^{l+1/2}+\lambda^{1/2}\right)+\varepsilon\cdot C_{2}\cdot\sqrt{\frac{d}{\lambda}}$ (48)
for every $\varepsilon>0$ and all $f\in H^{\sigma_{n,l}}(\Omega)$ with $\|f\|_{H^{\sigma_{n,l}}}\leq 1$ with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2C_{2}^{2}}\right)$ and constants $C_{1},C_{2}>0$ only depending on $\text{\sf{k}}_{n,l}$.
@@
The second claim combines the second part of Proposition 4.5 for the deterministic error bound with setting $C_{2}=\|\text{\sf{k}}_{n,l}\|_{\infty}^{1/2}$, obtaining (48). ∎